In the United States, 40% of the population has brown eyes. If 14 people are randomly selected, find the probability that at least 12 of them have brown eyes. Is it unusual to randomly select 14 people and find that at least 12 of them have brown eyes? Why or why not?
Answer #1

Answer

given that

n = 14

p = 0.40

and r = 12

we have to find the probability of at least 12 of them have brown eyes

using binomcdf

setting n = 14, p = 0.40 and k = 12

P(at least 12) = 1 -binomcdf(n,p,k-1)

= 1 - binomcdf(14,0.40,12-1)

= 1 - binomcdf(14,0.40,11)

= 1 - 0.9994

= 0.0006

= 0.0006*100 = 0.06%

So, this probability is unusual because it is less than 5% probability limit.